OpenStudy (anonymous):

does anyone know how to gernerate the equation of a circle with endpoints of diameter at (-3,5) and (7,11)

OpenStudy (anonymous):

ya. just use your equation of a circle

OpenStudy (anonymous):

ok i got (x-2)^2 + (y-8)^2= and then what?

OpenStudy (divanshu):

(x+3)(x-7) + (y-5)(y-11) = 0

OpenStudy (anonymous):

whoa thanks dude i thought u didn't like me lol

OpenStudy (anonymous):

how did you get that?

OpenStudy (divanshu):

sorry , if you felt i that

OpenStudy (anonymous):

haha its ok

OpenStudy (divanshu):

equation of any circle if the end points of its diameter are (a,b) and (p,q) is (x-a)(x-p) + (y-b)(y-q) =0

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

can you try to help with mine?

OpenStudy (anonymous):

sorry worm, i can't